A central angle with the measure 50° is drawn in a circle with circumference of 9π mi. What is the length of the intercepted arc to the nearest tenth of a mile

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

(50/360)*9pi miles

Answer by stanbon(75887)

: A central angle with the measure 50° is drawn in a circle with circumference of 9π mi. What is the length of the intercepted arc to the nearest tenth of a mile
------------------
P = 2pi*r
r = 9pi/2pi = 4.5 miles
----------------------
50 degrees = 50(pi/180) = 0.2778pi radians
------
Formula (arc length)/r = theta in radians
-----
Ans:: arc length = (4.5)*0.2778pi = 3.9270 miles
-----
Cheers,
Stan H.
------------